OpenStudy (anonymous):

find the value of x Select one: a. 6.9 b. 8 c. 19 d. 12

OpenStudy (anonymous):



OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

what is the 14 refering to? the whole base? Is it equilateral?

OpenStudy (aum):

Use the same technique as before. Label the vertices. Find similar triangles between the outermost triangle and the inner triangle whose dimensions are mostly known. Then solve for x.

OpenStudy (anonymous):



OpenStudy (anonymous):

@peterek0

OpenStudy (anonymous):

To be honest i have no clue. there are no specific angels nor information about other sides. Seems unsovable to me but i might be missing something obvious :)

OpenStudy (anonymous):

@aum : im getting 4.28 & thats not an answer choice...

OpenStudy (anonymous):

this is my work AD/BD=DB/CB x/6 = 6/14 x=60/14 = 30/7 = 4.28